### v2

# Statistical Principles for Platform Trials

Abstract
While within a clinical study there may be multiple doses and endpoints, across different studies each study will result in either an approval or a lack of approval of the drug compound studied. The False Approval Rate (FAR) is the proportion of drug compounds that lack efficacy incorrectly approved by regulators. (In the U.S., compounds that have efficacy and are approved are not involved in the FAR consideration, according to our reading of the relevant U.S. Congressional statute).
While Tukey’s (1953) Error Rate Familywise (ERFw) is meant to be applied within a clinical study, Tukey’s (1953) Error Rate per Family (ERpF), defined alongside ERFw, is meant to be applied across studies.
We show that controlling Error Rate Familywise (ERFw) within a clinical study at $5\%$ in turn controls Error Rate per Family (ERpF) across studies at 5-per-100, regardless of whether the studies are correlated or not. Further, we show that ongoing regulatory practice, the additive multiplicity adjustment method of controlling ERpF, is controlling False Approval Rate (FAR) exactly (not conservatively) at 5-per-100 (even for Platform trials).
In contrast, if a regulatory agency chooses to control the False Discovery Rate (FDR) across studies at $5\%$ instead, then this change in policy from ERpF control to FDR control will result in incorrectly approving drug compounds that lack efficacy at a rate higher than 5-per-100, because in essence it gives the industry additional rewards for successfully developing compounds that have efficacy and are approved. Seems to us the discussion of such a change in policy would be at a level higher than merely statistical, needing harmonizsation/harmonization. (In the U.S., policy is set by the Congress.)
Keywords: Platform trials; Error Rate per Family (ERpF); Error Rate Familywise (ERFw).
1 An motivating example: the HEALEY ALS Platform trial
Platform trials represent an emerging clinical trial design gaining popularity for their ability to conduct simultaneous testing of different treatments. More significantly, they enable the continuous incorporation of new treatments during the study, facilitating continuous learning. While platform trial design enhances resource efficiency and improves the likelihood of study participants receiving active treatments rather than placebos, the novelty of this design approach leaves ample room for exploration within the field of clinical trials, particularly regarding statistical principles due to the variety of potential outcomes.
While the excitement surrounding platform trials is undeniable, they do come with several limitations. For instance, platform trials are not universally applicable to all diseases, and meticulous planning is necessary to ensure their success. Nevertheless, despite these challenges, there have been instances of successful implementations of platform trials. One such case is the HEALEY ALS Platform trial, sponsored by Massachusetts General Hospital. Amyotrophic lateral sclerosis (ALS), commonly known as Lou Gehrig’s disease, is a progressive and fatal neuromuscular disease with the majority of ALS patients dying within 2 to 5 years of receiving a diagnosis.
[…]
In the HEALEY ALS Platform trial, the primary outcome is the ALS Functional Rating Scare-Revised (ALSFRS-R) score, which measures 12 aspects of physical functionality. It consists of 12 questions, each targeting a specific aspect of functional capability, with individual scores ranging from 4 (normal function) to 0 (complete loss of function). Thus, the aggregate score can span from a maximum of 48, indicating full functionality, to a minimum of 0, reflecting total loss of physical function. For a repeatedly measured continuous variable such as ALSFRS-R score, Mixed Models Repeated Measures (MMRM) analyses are typically employed.
A significant advantage of a Platform trial is the comparison of patients receiving each active treatment to placebo patients shared from other treatment arms (regimens). The sharing of placebo patients (which we will call Controls) from other treatment arms (regimens) allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, when employing models like MMRM to analyze repeatedly measured continuous variables such as the ALSFRS-R score, additional considerations emerge beyond the typical variability in estimated efficacy due to patient-to-patient differences within each treatment group. Concerns arise specifically regarding the influence of shared control samples. An overachieving shared control group could inadvertently suggest efficacy across all compounds in a platform trial, presenting challenges for sponsors. Conversely, an underperforming shared control group might lower the bar for all compounds to pass regulatory scrutiny too easily. This article addresses the necessity of planning for a sufficient total sample size to address these concerns while accommodating various randomization ratios that satisfy all stakeholders: patients, regulators, and sponsors.
2 False Approval Rate (FAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the purpose of Type I error rate control is to control the rate of incorrect regulatory decisions while balancing the benefits and risks to patients. When it is correctly done, Type I error rate control translates to False Approval Rate (FAR) control, i.e., controlling the rate of approving a treatment (e.g., drug) when it is not efficacious, across infinitely many studies (which are not identical, nor necessarily independent). In this paper, we refer to each clinical trail as a “study”.
2.1 Two layers of families and error rates
Within study
Within each study, there may be multiple doses and/or endpoints. The within-study family of inferential statements is the collection of (prespecified) statements to be made about them (such as confidence intervals for low dose efficacy in the primary endpoint and high dose efficacy in the secondary endpoint).
Across studies
Across all studies, the outcome of each study is whether the study compound is approved or not. The across-study family of inferential statements is the collection of statements to be made about whether the test compound of each study in the family has efficacy or not.
In general, when there are decisions to be made both within and across studies, multiple layers of families may exist. For instance, Ding et al., (2018) has a lower layer family comprising three confidence intervals for each single nucleotide polymorphism (SNP), assessing its efficacy across three possible genotypes. At a higher layer, there are confidence statements across all the SNPs in a genome-wide association study (GWAS), evaluating whether each SNP is linked to the phenotype, thus demonstrating the efficacy of a new treatment in comparison to a control. As another example, Han et al., (2022) has a lower layer family of three confidence intervals for each biomarker value, assessing the efficacy of a new treatment versus a control within three distinct groups: the marker-positive $g^{+}$ subgroup, marker-negative $g^{-}$ subgroup, the overall population $\{g^{+},g^{-}\}$. At a higher layer, there are confidence statements across uncountable many biomarker values.
In the Healey ALS platform trial, within each regimen, there exists a family of efficacy assessment on primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial forms a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
Principle 2.1 (Principle of layering of error rate controls).
When there are multiple layers of families of statistical inferences, error rate control can be applied in layers, to control the incorrect decision rate at each layer as well as for the overall decision-making process.
2.2 Tukey’s three Multiple Comparison error rates
It can perhaps be said that Tukey (1953), which has been reprinted as Tukey (1994), started Multiple Comparisons as a discipline in Statistics. The form of statistical inference that John W. Tukey preferred was confidence intervals. He defined the following error rates:
error rate per comparison
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of comparisons}}$
error rate per family
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of families}}$
error rate familywise
$=\frac{\mbox{number of erroneous families}}{\mbox{number of families}}$
In these definitions, each “statement” is a confidence interval, and an erroneous statement is a confidence interval that does not cover the true parameter value. An erroneous family is a family with at least one confidence interval not covering its true parameter value.
Note the now common phrasing familywise error rate transposes the error rate and familywise in Tukey’s original terminology. We believe the intent of the per in error rate per family was to have an indication of “how many?” after that. An error rate of 1 per a family of 20, or equivalently an error rate of 5 per a family of 100 (5-per-100 in short), both are $\mathbf{5\%}$ Tukey’s ordering of words is evocative because saying
  • error rate per a family of $k$ studies indicates control of errors is across studies
  • error rate familywise indicates control of errors is within each study
Therefore, instead of the transposed familywise error rate (abbreviated as FWER), we use Tukey’s original ordering error rate familywise (abbreviating it as ERFw). Also contrary to popular perception, it is not true that the only alternative to ERFw is false discovery rate (FDR). While FDR control may be appropriate for medical image reconstruction, it is well known that FDR control is inappropriate for controlling errors within a clinical trial because it can be manipulated by sponsors (see Section 6 of Finner and Roter 2001). Which multiple comparison error rate to control depends on the specific application as we will discuss below.
2.3 Error Rate per Family (ERpF) control guards regulatory False Approval Rate (FAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to incorrect claim of additional efficacy on the drug label. Since any incorrect statement can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise within each study since the early 2000s. We will show that across studies, controlling error rate per family appropriately balances the risk and benefit to patients, regardless of whether the studies are independent or correlated in families of Platform trial studies.
Even though individual studies and those in platform trials are conducted at overlapping times, we can think of them as occurring in a (countably) infinite sequence of clinical trials, indexed as Study $\ldots,-1,0,1,\ldots$. Consider a finite number of $k$ studies, some of which are in platform trials while others are not. Within these $k$ studies, some compounds have efficacy while others do not. Studies with efficacious compounds that fail, failing to reject the null hypothesis of no-efficacy, do not lead to incorrect regulatory approval. We thus isolate those studies with compounds that lack efficacy (and therefore should not gain regulatory approval) as studies with index in $N$, and let $n$ denote its cardinality (the number of studies with index in $N$). To control the error rate per family (ERpF) of studies in $N$ at $5\%\times n$ per a family of $n$ studies means, as $n\rightarrow\infty$, the proportion of compounds lacking efficacy that incorrectly gain regulatory approval approaches $5\%$.
Different treatment arms within a Platform trial may share a common control arm and therefore not independent. Among the infinitely many studies across independent as well as those in Platform trials, the following theorem allows us to show if error rate familywise (ERFw) of every study within each Platform trial is set at the 5% level, this dependence does not affect regulatory false approval rate being no more than 5-per-100 (or 5%), so long as dependence is short term (within each Platform trial only, while studies in different Platform trials remain independent).
Theorem 2.1 (Law of Large Numbers for random variables with short term dependence).
Let $X_{l},X_{2},\ldots,$ have means $\mu_{1},\mu_{2},\ldots$ variances $\sigma^{2}_{1},\sigma^{2}_{2},\ldots$ and covariance $Cov\{X_{i},X_{j}\}$ satisfying
$Cov\{X_{i},X_{j}\}\leq\rho_{j-i}\sigma_{i}\sigma_{j}~{}(i\leq j)$ (2.1)
where $0\leq\rho_{m}\leq 1$ for all $m=0,1,\ldots$. If the series $\sum_{1}^{\infty}\rho_{m}$ and $\sum_{1}^{\infty}\sigma^{2}_{i}(\log i)^{2}/i^{2}$ are both convergent, then as $n\rightarrow\infty$
${n}\sum_{i=1}^{n}X_{i}-\frac{1}{n}\sum_{i=1}^{n}\mu_{i}\rightarrow 0~{}{w.p.1}.$1 . (2.2)
Proof.
See Theorem E on page 27 of Serfling, (1980). ∎
Let $\ldots,Y_{-1},Y_{0},Y_{1},\ldots$ be the indicator functions of past, present, and future studies that lead to efficacy claims from rejection of the null hypothesis that the compound studied lacks efficacy. Among studies with index in $N$ that should not lead to regulatory approvals, we relabel their $Y_{i}$ as $X_{i}$. We show that controlling ERFw within each study controls regulatory false approval rate ERpF across studies, regardless of whether these studies are in platform trials or not.
Let $X_{i}$ be the indicator of a Type I error being committed as a result of Study $i,~{}i\in N,$ a study with compound that lacks efficacy,
$X_{i}=\left\{\begin{array}[]{ll}1&\mbox{if a Type I error is made as a result of Study}~{}i,i\in N,\\ 0&\mbox{otherwise}.\end{array}\right.$ (2.3)
then $\mu_{i}\leq 0.05$ for all $i\in N$. For $X_{i},X_{j}$ within the same Platform trial, the correlation $\rho_{j-i}$ in (2.1) is between zero and one. But with short term dependence, for $X_{i},X_{j}$ in different Platform trials, $\rho_{j-i}=0$. For example, if the practical maximum number of study arms within a single Platform trial is 100, then the $\rho_{j-i}$ in (2.1) would satisfy
$\rho_{j-i}\left\{\begin{array}[]{ll}\leq 1&\mbox{if }|j-i|\leq 100\\ =0&\mbox{if }|j-i|>100.\end{array}\right.$
Let $m$ = $|j-i|$, then $\rho_{m}=0,~{}m=101,102,\ldots$ for Platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series. As $\frac{1}{n}\sum_{i=1}^{n}\mu_{i}\leq\frac{1}{n}\sum_{i=1}^{n}0.05=0.05$, the long run regulatory FAR will not exceed 5-per-100 ($5\%$) if the Type I error rate familywise (ERFw) of each study is controlled at $5\%$. Consequently, for a regulatory agency such as the Food and Drug Administration (FDA) or the European Medical Agency (EMA) to not approve medicine that are ineffective more than 5-per-100 across infinitely many studies, controlling the error rate familywise (ERFw) of each study at the 5% level guarantees that, according to Theorem 2.1.
2.4 Additive multiplicity adjustment for regulatory approval
Congress of the United States enacted in 1938 the Food, Drug, & Cosmetic Act. Together with its 1962 Kefauver-Harris Amendment, they require pharmaceutical companies to submit evidence of new drugs’ safety and effectiveness for approval to market to the public. To balance risk-benefit to the public, the U.S. Food and Drug Administration (FDA) has been allowing five incorrect approvals per 100 ineffective compounds, 5-per-100, as follows.
The FDA requires control of the error rate familywise (ERFw) within each study to be at $5\%$ strongly, so that even for a complex study (with multiple doses and endpoints in combination), the chance that an ineffective compound will show efficacy is no more than $5\%$. Then, the regulatory mission is accomplished across any family of $n$ studies with inefficacious compounds because the error rate per family (ERpF) is controlled at 5-per-100 ($5\%\times n$), as follows.
For a family of $n$ studies of inefficacious compounds with index in $N$, let $V$ be the number of studies with Type I errors in this family. Then $V=\sum_{i\in N}X_{i}$ where $X_{i}$ is as in (2.3) the indicator that a Type I error is committed as a result of Study $i,~{}i\in N$. Expectation of a sum of random variables simply add. Suppose $P\{X_{i}=1\}=5\%$ exactly (not conservatively) for all $i,~{}i\in N$. Then
$E[V]$ $=$ $\sum_{i\in N}P(X_{i}=1)=5\%\times n$
exactly, not conservatively, regardless of whether $X_{i},~{}i\in N,$ are independent, positively dependent, or negatively dependent. So in terms of the expected number of incorrect regulatory approvals, controlling the Type I error rate familywise (ERFw) of each study exactly at 5% leads to controlling the incorrect regulatory approval rate at exactly 5% for a family of $n$ studies with inefficacious compounds (which we phrase as 5-per-100). This is as expected of a regulatory agency, balancing risk and benefit to patients, contrary to the perception given by Bai et al (2020).
3 False Discovery Rate (FDR) in a regulatory setting
3.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed for Type I error rate control across studies in Platform trials (add reference here). FDR is the expected proportion of rejections (among rejections both correct and incorrect) that are incorrect [^2] Rejecting a false null hypothesis is “correct”, while rejecting a true null hypothesis is “incorrect”.. Let $H_{0i},i=1,\ldots,k$ be the null hypotheses that each compound in a Platform trial with $k$ studies lacks efficacy, then
$\mathrm{FDR}$ $=$ $\mathrm{E}\left[\frac{\mbox{number of rejected }H_{0i}\mbox{ thatare true}}{\mbox{total number of rejected }H_{0i}}\right]$
$=$ $\mathrm{E}\left[\frac{\mbox{number of inefficacious compounds incorrectly approved }}{\mbox{total number of compounds approved, both efficacious and inefficacious}}\right]$
While current regulatory practice of controlling FAR (which is Tukey’s per family error rate across studies) limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100.
Suppose a Platform trial has $n^{\mathrm{true}}$ true no-efficacy nulls $H_{0i}^{{\scriptsize\mbox{true}}},i=1,\ldots,n^{\mathrm{true}}$, the expected number of incorrect rejections (incorrect approvals) is just
$\mathrm{ERpF}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ (3.2)
Shouldn’t the $v$ here be $i$? So to compute $E[V]$ we add up the Type I error rate of each study. If FDR were fixed, then to do this calculation we can try reverse-engineering $P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ and adding them up. For easier subsequent comparison with FDR (which is a fraction), we can divide ERpF by the total number of drug compounds that should not be approved
$\dfrac{\mathrm{EFpF}}{n^{\mathrm{true}}}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{n^{\mathrm{true}}}P\{V=v\}$ (3.3)
Now suppose a Platform trial has $n^{\mathrm{false}}$ extremely false no-efficacy nulls (that will almost surely get rejected), then
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{v+n^{\mathrm{false}}}P\{V=v\}$ (3.4)
$\leq$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{0+n^{\mathrm{false}}}P\{V=v\}.$ (3.5)
How ERpF and FDR count the number of incorrect approvals $v$ differ is evident by comparing (3.4) with (3.3). FDR attenuates the number of incorrect rejection $v$ by the number of correct rejection of false nulls.
From (3.5), we see $n^{\mathrm{false}}$ being not small (relative to $v$) allows inflation of $P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ while keeping FDR small. For simplicity, suppose each Platform trial has $k=2$ studies only, and suppose the compound in Study $w$ is highly efficacious but the compound in Study $v$ is in-efficacious:
  • $H_{0w}^{{\scriptsize\mbox{false}}}$ is so false that $P\left\{\mbox{Reject }H_{0w}^{{\scriptsize\mbox{false}}}\right\}=1$;
  • $H_{0v}^{{\scriptsize\mbox{true}}}$ is true so rejecting it constitutes the only incorrect rejection.
In this scenario, $H_{0v}^{{\scriptsize\mbox{true}}}$ can be tested at a Type I error rate (ERFw) of $10\%$ (within Study $v$) while keeping FDR in this Platform trial of a family of 2 studies at $5\%$, because
$FDR$ $=$ $E\left[\frac{V}{R}\right]$
$=$ $\left(\dfrac{1}{1+1}\right)\times P\left\{\mbox{Reject }H_{0v}^{{\scriptsize\mbox{true}}}\right\}$
$=$ $\frac{1}{2}\times 10\%=5\%$
So, if there is a sequence of many Platform trials with $k=2$ studies following this scenario, then by the LLN (2.1), the False Approval Rate (FAR) is 10-per-100 (not 5-per-100).
Switching to FDR control would thus be a policy change: For every efficacious compound approved, regulators reward the industry with an additional fungible False Approval Rate (FAR) token [^3] We term the tokens fungible because any company can cash them in for any compound. toward approval of inefficacious compounds. We raise this caution as we do not quite see in the Food, Drug, & Cosmetic Act or the Kefauver-Harris Amendment language indicating intent that guarding against incorrectly approving inefficacious drug compounds should be affected by the number of efficacious compounds correctly approved.
3.2 Integrity of error rates against manipulation is desirable
Whereas the False Approval Rate (FAR) is only concerned with the rate at which inefficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter (2001), we wonder whether that might subject FDR to manipulation as follows.
If FDR were to be controlled across studies in Platform trials involving multiple drug companies, seems to us an astute company would not put forth their best compound, leaving that to other companies. Even in Platform trials involving a single sponsor, we wonder whether there would be temptation to include highly efficacious compounds of low commercial value, to benefit compounds with low efficacy but high commercial value.
4 Stability of FAR and Conditional Assessment
While reassuring that a 5% ERFw within each study results in an ERpF of 5% across all the studies, there is a concern regarding the stability of ERpF across Platform trials.
4.1 Conditional Stability of FAR is desirable
Section F titled Multiplicity of FDA’s draft guidance on Master Protocols states
the probability distribution for the number of type I errors should be considered both in evaluating a proposed design and analysis plan and in evaluating the persuasiveness of results.
An error rates which is an expectation is just the expected value aspect of the probability distribution of errors made. For any unconditional error rate, even with it being controlled at 5%, it could be the case that some families of studies (with particular characteristics of their studies within) can have actual error rates that are rather variable (varying from 1% to 9% say), while other families of studies can have error rates that are more stable (ranging from 4% to 6% only say). So, in addition to the unconditional (expected) error rate, variability of the actual error rate should be considered as well, conditionally given characteristics of studies within families. The specific characteristics of concerns from sharing a control in a platform trial are as follows:
Mean $\hat{\mu}_{SC}$
An under-performing or over-achieving estimated mean $\hat{\mu}_{SC}$ of the shared control makes getting approval currently easier or harder for all the new compounds, with under-performing $\hat{\mu}_{SC}$ especially of concern to regulatory agencies;
Correlation $\hat{\rho}_{SC}$
A high realized (overall) correlation measure $\hat{\rho}_{SC}$ among the test statistics associated with the new compounds can substantially inflate the variability of $V$, the number of incorrect discoveries/approvals. Note that correlation this can be represented as a product of a function of the ratio of sample size $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$. Specifically, $\hat{\rho}_{SC}=\lambda_{i}\lambda_{j}$ for all $1<=i<j<k$. These $\lambda$ values can easily be found via existing software. We favor this notation due to the ease of extension to more complicated structures. For the HEALEY ALS pltform tril, $\lambda=$ (0.674, 0.722, 0.671, 0.690) to a close approximation.
To facilitate the assessment of impact of correlations, analogous to commerce on the internet keeping track of cookies, we propose an informational “cookie” for each treatment/regimen of a platform trial to be kept track of.
Control Sharing Matrix
Cookie for each treatment/regimen should include the exact time interval of the enrollment period of each regimen, so that by knowing the “overlap” for the enrollment periods of a pair of regimens, one can compute the pairwise covariance and correlation for the active treatment versus Control comparisons.
$n$ A B C D
A 122 122 164 103
B 122 120 122 103
C 164 122 120 103
D 103 103 103 75
Table 1: Control Sharing Matrix (CSM) for the HEALEY ALS Platform trial (inferred as best we can from Results posted at clinicaltrials.gov): For each regimen, the corresponding diagonal element is the number of patients randomized to the active treatment. For each pair of regimens, the corresponding off-diagonal element is the number of patients that they have in common in the shared Control that each active treatment is compared to.
A B C D
A 1 0.498 0.425 0.478
B 0.498 1 0.496 0.476
C 0.425 0.496 1 0.476
D 0.478 0.476 0.476 1
Table 2: Matrix of correlations between $\hat{\mu}_{i}-\hat{\mu}_{SC_{ij}}$ and $\hat{\mu}_{j}-\hat{\mu}_{SC_{ij}}$
Within each new treatment arm, randomization ratio $n_{Rx}:n_{C}$ of new compound to the shared control can range from a low of $1:1$ (in which case the sample size ratio of the shared control to each new compound becomes $k:1$) to a high of $k:1$ (in which case the sample size ratio of the shared control to each new compound becomes $1:1$). Given a fixed total sample size for the platform trial, choice of the randomization ratio $n_{Rx}:n_{C}$ involves multiple considerations.
Patients would prefer a higher $n_{Rx}:n_{C}$ randomization ratio since higher it is higher the chance of receiving a new treatment compound.
From an individual drug sponsor’s point of view, a benefit of joining a platform trial is pooling control patients allows a randomization ratio higher than the $1:1$ in a standalone study, to more accurately estimate effects of the new compounds. However, that is balanced by the master protocol sponsor’s view that lower the $n_{Rx}:n_{C}$ randomization ratio higher the sample size of the shared control, which lessens the chance that an overachieving shared control sample causes all the new compounds difficulty demonstrating efficacy.
On the other hand, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable.
For example, none of the four drugs tested in Regimens ABCD of the HEALEY ALS Platform trial demonstrated efficacy in the primary endpoint ALSFRS-R, [^4] Active treatment for Regimen A was Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine. which can be due to either the compounds lacking efficacy, or from an overachieving shared control, or both.
Choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these considerations into account. Our app lets one find the sweet spot.
4.2 Effect of mean and correlation induced by sharing control on $V^{\star}$
It is to be noted that while both mean and correlation induced by sharing control on $V^{\star}$ individually affect the distribution of $V$ and thus $V^{\star}$, we need to consider the simultaneous effect on the distribution. The variance of $V$ in particular is of note. As seen in Figure 1, for cases when $\hat{\mu}_{SC}$ is greater than the treatment arm mean $\hat{\mu}_{i}$, the performance of the standard deviation of $V$ is much more stable. On the other hand, when $\hat{\mu}_{SC}$ is less than $\hat{\mu}_{i}$, the behavior of the standard deviation is much more volatile. This tells that in cases like this, it is beneficial to look at the whole behavior of the probability distribution of V rather than single instances.
Figure 1: The behavior of standard deviation for V is heavily affected by correlation and sample control. We assume the true $\mu_{sc}=0$, i.e. it is non efficacious. As we decrease the value of our $\hat{\mu}_{SC}$, the standard deviation of $V$ shows less stability.
$\hat{\mu}_{SC}$ = -3
$\hat{\mu}_{SC}$ = -1.5
$\hat{\mu}_{SC}$ = 0
$\hat{\mu}_{SC}$ = 1.5
$\hat{\mu}_{SC}$ = 3
Figure 2: Snapshot of different standardized $\hat{\mu}_{SC}$ values on correlation’s effect on standard deviation where $\mu_{i}=0$. As can be seen, when the standardized $\hat{\mu}_{SC}$ is less than the treatment mean, the behavior is less simple whereas for $\hat{\mu}_{SC}$ is greater than the treatment mean, the standard deviation of V decreases as correlation increases.
4.3 Effect of correlation induced by sharing control on $V^{\star}$
Let us say there are 20 new treatments in a platform trial, and 10 of the treatments show efficacy. Let $V^{\star}$ denote the number of treatments out of those ten with their confidence intervals for efficacy not covering their true parameter values, then $E[V^{\star}]=0.05\times 10=0.5$. If these new treatment arms were independent, then $StdDev(V^{\star})$, the standard deviation of $V^{\star}$ (under the null) is 0.69 which is larger than the mean $E[V^{\star}]=0.5$. Further, correlation among the test statistics associated with the new compounds (induced by sharing a control) makes the variability of $V^{\star}$ (the number of incorrect approvals) even bigger. In the scenario above, a common correlation of 0.5 among the test statistics increases $StdDev(V^{\star})$ from 0.69 to to 1.16. This ever-present variability of the actual error rate in a particular family of clinical trials is a fact of life that we believe should receive more attention in practice, regardless of the error rate being ERpF or ERFw. We show how this calculation is done below.
Let $N$ denote the subset of index $\{1,\ldots,k\}$ of compounds that lack efficacy. For $i\in N$, let $Z_{i}$ be the indicator that the $i^{th}$ confidence interval erroneously show efficacy. Then $V^{\star}=\sum_{i\in N}Z_{i}$. Since $Var(V^{\star})=\sum_{i}Var(Z_{i})+\sum_{i\neq j}Cov(Z_{i},Z_{j})$ and $Var(Z_{i})=\alpha(1-\alpha)$, we just need the covariances $Cov(Z_{i},Z_{j}),i\neq j,$ to compute $Var(V^{\star})$.
There are various ways of computing the non-coverage events covariances among the treatment arms in a platform trial. For proof of concept, consider platform trials in therapeutic areas where it is reasonable to take each pair of $(T^{\star}_{i},T^{\star}_{j}),i\neq j,$ as bivariate standard Normal. Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM) are examples of such therapeutic areas. Then
$Cov(Z_{i},Z_{j})=P\{T^{\star}_{i}>z_{\alpha}\cap T^{\star}_{j}>z_{\alpha}\}-\alpha^{2},$ (4.1)
which depends on $\rho_{ij}=Corr(T^{\star}_{i},T^{\star}_{j})=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$.
Each subset $N$ has its own $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio between sample sizes in the arms of treatments with index in $N$ and in the shared control arm but also on other covariates such as the baseline measurement of each patient. In assessing stability of error rate control of compounds with apparent efficacy, we consider the observed treatment outcomes as well as the covariates as known, So we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ for the $N$ of compounds that show efficacy. For modeling-based analyses, such computations are possible.
A B C D
A 0.0475 0.0093 0.0083 0.0087
B 0.0093 0.0475 0.0092 0.0096
C 0.0083 0.0092 0.0475 0.0086
D 0.0087 0.0096 0.0086 0.0475
Table 3: Matrix of variances of $Z_{i}$ (diagonal) and covariances between $Z_{i}$ and $Z_{j}$ (off-diagonal) with ERpF set at $5\%\times 4=0.2$ for Regimens ABCD of the ALS Platform trial. The elements sum to 0.2974 for $Var(V^{\star})$, so $StdDev(V^{\star})$ = 0.5454, which is larger than the ERpF of 0.2.
Table 4 shows $StdDev(V^{\star})$ the variability of $V^{\star}$ increases with increasing $\rho_{ij}$.
$k$ $E(V^{\star}$ ) $StdDev(V^{\star})$
Unconditional Independent studies Hypothetical Platform $k:1$ randomization
$n^{\mathrm{true}}$ ERpF $\rho_{ij}=0$ $\rho_{ij}=0.3$ $\rho_{ij}=0.5$
5 0.25 0.49 0.57 0.66
10 0.5 0.69 0.94 1.16
20 1 0.97 1.65 2.15
40 2 1.38 3.02 4.12
Table 4: For a family of studies, if $StdDev(V^{\star})>E(V^{\star})$, that is, standard deviation of the number of errors is larger than the mean of the number of errors, then one might consider stability of the ERpF control to be of concern. In a Platform trial, pairs of $T^{\star}_{i}$ and $T^{\star}_{j}$ are correlated due to sharing of Control patients. Assuming $\rho_{ij}$ is common for all pairs, variability of $V^{\star}$ increases with increasing $\rho_{ij}$. For illustration, consider a Platform trial with $k$ new compounds using a $k:1$ compound-to-placebo randomization (which makes the sample size for comparing each new compound with the shared Control $1:1$). With ERpF set at $5\%\times n^{\mathrm{true}}$ where $n^{\mathrm{true}}$ is the number of in-efficacious compounds, we see independent studies reach stability by family size 20, but that is not the case for studies in Platform trials. Stability of ERpF control is thus more of a concern in Platform trials.
4.4 Conditional assessment of FAR control
A foundational principle in (frequentist) Statistics is
Principle 4.1 (The Conditionality Principle).
If, among many experiments that can inform on an unknown parameter, one is randomly selected and performed, then the information available about the unknown parameter depends only on the experiment actually performed.
See J. O. Berger (1985) page 31 and Casella and R. L. Berger (1990) page 268 for details.
To motivate the application of this principle to conditional confidence, consider Fieller’s confidence interval for the ratio of the means of two Normal distributions (Fieller 1954) which unconditionally (over infinitely many replications of a study) has confidence level exactly equal to $(1-\alpha)$. Interestingly, depending on the data observed in each particular study, Fieller’s confidence interval can be an interval, or the entire real line, the complement of an interval, or the empty set (see Casella and Berger 1990, page 459). To us, conditional on Fieller’s interval being the entire real line for a particular study, a conditional confidence of 100% should be stated for that study. Likewise, conditional on Fieller’s interval being the empty set for a particular study, a conditional confidence of 0% should be stated for that study, not averaging the coverage probability over studies that have not been performed that would result in Fieller’s confidence interval being an interval, the complement of one, or the entire real line. In this spirit, while taking comfort that a 5% ERFw within each study results in an unconditional ERpF of 5% across all the studies, we recommend assessing the conditional ERpF within each Platform trial conditional on the observed shared Control outcome, to be sure they do not deviate too much from the unconditional ERpF.
In our Platform trials setting, let $T_{i},i=1,\ldots,k,$ be the statistics for comparing the $k$ new compounds against a shared control in a platform trial. For proof of concept, consider platform trials in therapeutic areas where $T_{i}$ can be pivoted to form a $100(1-\alpha)\%$ confidence interval for the primary efficacy endpoints44footnotetext: which excludes the score test version of the logRank test, so that
$T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$ (4.2)
the pivoted $T_{i}$ has a null distribution free of the efficacy parameter. Neurodegenerative diseases (such as Alzheimer’s) and Type 2 Diabetes Mellitus (T2DM) are examples of such therapeutic areas. For simplicity of illustration, suppose $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are Normally distributed and the sample sizes are large enough so that $SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}$ in (4.2) can be considered essentially a constant.
Let $V^{\star}$ denote the number of new treatments in a platform trial with their confidence intervals for efficacy not covering their true parameter values, Then sample mean $\hat{\mu}_{SC}$ of the shared control under-performing or over-achieving affects the distribution of $V^{\star}$. Correlation $\hat{\rho}_{SC}$ among the test statistics associated with the new treatments (induced by sharing a control) affect the distribution of $V^{\star}$ as well.
4.4.1 Effect of sample mean of shared control on $V^{\star}$
(a) $\hat{\mu}_{SC}=0.257$
1 SE above $\mu_{SC}$
(b) $\hat{\mu}_{SC}=-0.257$
0.5 SE below $\mu_{SC}$
(c) $\hat{\mu}_{SC}=-0.514$
1 SE below $\mu_{SC}$
Figure 3: Probability distribution of $V$, the number of times a type I error is committed. We assume 4 treatment arms with an $\alpha$ level of 0.05, giving an ERpF of $0.05\times 5=0.25$, a total sample size of 800, a ratio of 4:1 for treatment arm to control sample size allocation gives an unconditional correlation of new treatment arm minus control estimates of 0.5. We additionally assume that the true shared control mean $\mu_{SC}$ and treatment arm means $\mu_{i}$ are equal to 0 (but this can easily be modified to other numbers). In Figure 3(a), the distribution of V is fairly similar to the baseline. Since the $\hat{\mu}_{SC}$ is higher, treatment arms have more difficulty gaining approval. On the other hand, in Figure 3(b) and 3(c), since our $\hat{\mu}_{SC}$ value is much lower than out $\mu_{i}$, erroneous approvals become much easier, and thus we have higher weight on higher values of $V$.
We consider 2 cases: (1) positive $\hat{\mu}_{SC}$ and (2) negative $\hat{\mu}_{SC}$. For sake of simplicity, we consider the cases where all treatment arms are non-efficacious, i.e. $\mu_{SC}=0$.
For cases with a positive $\hat{\mu}_{SC}$, the situation is of concern for sponsors, who want treatments to be approved. The more positive the $\hat{\mu}_{SC}$ treatment arms have a more difficult time getting approved, resulting in $V$ having more weight on lower numbers of erroneous approvals and thus translating to having a higher $V^{\star}$. For regulators, especially in cases where treatment arms are non-efficacious, however, this does not pose an issue as the chances of erroneously approving a treatment are low as can be as seen in Figure 3(a).
For the cases with negative $\hat{\mu}_{SC}$, the situation becomes of concern to regulators. It is much easier to approve treatment arms. Due to this, however, the issue of erroneous approvals becomes much more apparent. In fact, it could even result in almost all the treatment arms being erroneously approved if the $\hat{\mu}_{SC}$ was low enough. While pharmaceutical companies may desire this outcome, this is not desirable from a consumer or regulation standpoint as it leads to non-efficacious drugs getting erroneously approved (and thus increase $V^{\star}$).
Let us assume a simple case where we have 5 trials that are not efficacious. We adjust $\hat{\mu}_{SC}$ so that it is standardized by $SE_{SC}$. If we have a high $\hat{\mu}_{SC}$ 0 such that all treatment arms would not be approved, our $V^{\star}$ would equal 5. On the other hand, we consider a $\hat{\mu}_{SC}<0$ such that all treatment arms would be erroneously approved. This would result in a $V^{\star}$ of 5.
This problem can be built upon to address different scenarios to account for the sum of the treatment arm sample size $n_{SC}$ and control arm $i$ sample size $n_{i}$ and correlation $\rho=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$. This allows for consideration of $SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}$ as well, which is dependent on the population standard deviation $\sigma$ and the sample sizes $n_{SC}$ and $n_{i}$.
In general, having a larger shared control sample mean $\hat{\mu}_{SC}$ results in new treatment arms having a more difficult time getting approved, thus resulting in a lower $V^{\star}$. This makes intuitive sense since our comparison of control for efficacy would be higher so we would need a higher treatment arm score to gain approval. In the same manner, having a lower $\hat{\mu}_{SC}$ results in easier approval, resulting in more confidence intervals for efficacy not covering the true parameter values (and thus a higher $V^{\star}$ value). This effect on $V^{\star}$ demonstrate the importance of its consideration when testing for efficacy.
5 A Confident Platform Random Sizer App
An app has been created to not only help demonstrate the effect of conditioning on $\hat{\mu}_{SC}$, but also provide information about the calculated unconditional correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$ and . We use the HEALY ALS trial as an example case. While all four drugs tested in Regimens ABCD’s lack of efficacy could be due to either the compound lacking efficacy or an overachieving shared control, for the sake of this example, we will assume both. Each arm had approximately 120 patients in the treatment and approximately 40 in the control, maintaining a treatment-to-control sample size ratio of 3:1. We can set the app so the maximum number of treatments to be erroneously approved is 4 with a 120:40 (3:1) sample ratio. Based on this information, we can then input values into our sliders. Note that ALS has a population standard deviation of 6.5 (i.e. $\sigma=6.5)$ which is used to calculate the standard errors internally in the app (See App: User Slider Controls)
Figure 4: Although none of drugs in Regimens ABCD in HEALY ALS trial were determined as efficacious, if we had an underperforming shared control sample mean, i.e. $\hat{\mu}_{SC}<0$, there would be a chance drugs could be approved with ease as shown above with a standardized $\hat{\mu}_{SC}=-1.5$
App: User Slider Controls $\alpha$ for Each Rx Arm: Type I error rate for each treatment arm. Note that ERpF is $k\alpha$
Assume $\alpha=0.5$ for each treatment arm.
Standardized $\hat{\mu}_{SC}$: The shared control sample mean $\hat{\mu}_{SC}$, standardized by the standard error of the shared control sample mean
We assume an overachieving $\hat{\mu}_{SC}$, i.e. $\hat{\mu}_{SC}>0$ Different shared control values would affect the number of erroneous approvals
Maximum Number of Rx Arms Erroneously Approved: $n$, the number of treatment arms approved when they are non-efficacious. Assumes $n=k$, i.e. all treatments are non-efficacious.
$k=4$ so set $n=4$
Total Sample Size: Sum of sample sizes from all treatment arms and the shared control
[…]
[Uncaptioned image] Unconditional Correlation Between New Treatment and Control Estimates: correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
While not given, this information can be found via the sample size ratio below.
Sample Size Ratio: Treatment to control sample size ratio for sample allocation. Note that while the app allows for one to go higher, it is generally not advised to go higher than $k:1$.
textit3:1 treatment-to-control sample size ratio
App: Output Distribution of V: Gives the expected value, standard deviation, and the distribution of $V$, the number of treatment arms erroneously approved.
Since we assume an overachieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
Sample Allocation: Gives the recommended sample sizes for each treatment arm and the shared control (rounded to nearest whole number)
The study utilizes as 120:40 ratio, so a ratio of 3. The allocation approximately matches that in the study
[Uncaptioned image] Standard Errors: Gives standard errors for estimated control mean, treatment arm means, and the mean difference between control and treatment arm
Standard errors were calculated based on the standard deviation $\sigma$ = 6.5
By choosing the value for the sample size ratio, the unconditional correlation of the new treatment arm and control estimates gets calculated automatically for the user, as is the case for the ALS study. This is beneficial to regulators in particular, who would have more interest invested in controlling this to stabilize the distribution of the control sample mean. By having lower correlation (and thus lower randomization ratio), $SE_{\hat{\mu}_{SC}}$ is lowered as well, making the estimate for less variable.
While this example utilizes an existing study, this application could also demonstrate its usage when designing a study, specifically into considerations of stability of the estimates of shared control and (unconditional) correlation between the difference between $T_{i}^{\star}$ and $T_{j}^{\star}$. In particular, if regulators demand a certain correlation threshold to be met, sponsors are able to determine the highest sample ratio randomization that can achieve said threshold. The app can be used to find the sweet spot where all parties of interest can be satisfied.
6 Conclusion
In this paper, we examined the error rates within clinical studies, particularly focusing on platform trials. Our findings offer significant insights into the control of error rates and their implications across various study designs, especially in the context of regulatory and clinical settings.
We demonstrated that controlling the ERFw within a single study at a threshold of $5\%$ effectively controls the ERpF across multiple studies at a rate of 5-per-100. This finding holds irrespective of the correlation between studies, including those grouped within platform trials. Furthermore, our analysis supports that current regulatory practices, which involve controlling the ERFw within a study, limit the False Approval Rate (FAR) to 5-per-100. This approach ensures that no more than five ineffective drug compounds receive approval per hundred evaluated, thus safeguarding public health while fostering innovation.
A significant advantage of a Platform trial is that multiple treatments are evaluated simultaneously against a shared control, which allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, the stability of regulatory FAR across Platform trials is of serious concern since shared controls introduce dependencies among treatment evaluations and therefore increases the variability of ERpf. As we demonstrated in our analysis, dependencies among treatment evaluations increases with the treatment/control randomization ratio. From patients perspective, a higher treatment/control randomization ratio will allow higher chance of receiving a new treatment compound. From an individual drug sponsor’s point of view, a higher treatment/control randomization ratio through participating platform trial allows more accurate estimate on the effects of new compounds. However, from the master protocol sponsor’s point of view, lower the the treatment/control randomization ratio allows increasing the sample size of the shared control, and therefore lessens the chance that an overachieving shared control sample causes all the new compounds have difficulty to demonstrate efficacy. Additionally, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable. To balance the different perspective from different stake holder, our Confident RandomSizer app offers a tool to identify optimal treatment/control randomization ratio given a fixed sample size.
In conclusion, maintaining stringent control of error rates within clinical trials is crucial for ensuring the reliability of study outcomes and the safety of approved treatments. As clinical trial designs become increasingly complex, particularly with the rise of platform trials, the statistical community and regulatory bodies must continuously adapt and refine their methodologies to address these challenges effectively
7 Glossary
7.1 Terms
  • ERpF: Error rate per family
  • FAR: False Approval Rate
  • FDR: False Discovery Rate
  • FWER: Family wise error rate
7.2 Notation
  • $\mu_{SC}$ Shared control sample mean
  • $\hat{\mu}_{SC}$ Shared control sample mean estimate
  • $V^{\star}$