### v3

# Statistical Principles for Platform Trials

Abstract
While within a clinical study there may be multiple doses and endpoints, across different studies each study will result in either an approval or a lack of approval of the drug compound studied. The term False Approval Rate (FAR) is the term this paper utilizes to represent the proportion of drug compounds that lack efficacy incorrectly approved by regulators. (In the U.S., compounds that have efficacy and are approved are not involved in the FAR consideration, according to our reading of the relevant U.S. Congressional statute).
While Tukey’s (1953) Error Rate Familywise (ERFw) is meant to be applied within a clinical study, Tukey’s (1953) Error Rate per Family (ERpF), defined alongside ERFw, is meant to be applied across studies.
We show that controlling Error Rate Familywise (ERFw) within a clinical study at $5\%$ in turn controls Error Rate per Family (ERpF) across studies at 5-per-100, regardless of whether the studies are correlated or not. Further, we show that ongoing regulatory practice, the additive multiplicity adjustment method of controlling ERpF, is controlling False Approval Rate (FAR) exactly (not conservatively) at 5-per-100 (even for Platform trials).
In contrast, if a regulatory agency chooses to control the False Discovery Rate (FDR) across studies at $5\%$ instead, then this change in policy from ERpF control to FDR control will result in incorrectly approving drug compounds that lack efficacy at a rate higher than 5-per-100, because in essence it gives the industry additional rewards for successfully developing compounds that have efficacy and are approved. Seems to us the discussion of such a change in policy would be at a level higher than merely statistical, needing harmonizsation/harmonization. (In the U.S., policy is set by the Congress.)
Keywords: Platform trials; Error Rate per Family (ERpF); Error Rate Familywise (ERFw).
1 An motivating example: the HEALEY ALS Platform Trial
Platform trials represent an emerging clinical trial design gaining popularity for their ability to conduct simultaneous testing of different treatments. More significantly, they enable the continuous incorporation of new treatments during the study, facilitating continuous learning. While platform trial design enhances resource efficiency and improves the likelihood of study participants receiving active treatments rather than placebos, the novelty of this design approach leaves ample room for exploration within the field of clinical trials, particularly regarding statistical principles due to the variety of potential outcomes (Koenig et al.,, 2023).
While the excitement surrounding platform trials is undeniable, they do come with several limitations. For instance, platform trials are not universally applicable to all diseases. Meticulous planning is also necessary to ensure their success (Koenig et al.,, 2023).. Nevertheless, despite these challenges, there have been instances of successful implementations of platform trials. One such case is the HEALEY ALS Platform Trial, sponsored by Massachusetts General Hospital.
[…]
ALS is a progressive and fatal neuromuscular disease with the majority of ALS patients dying within 2 to 5 years of receiving a diagnosis. In the HEALEY ALS Platform Trial, the primary outcome is the ALS Functional Rating Scare-Revised (ALSFRS-R) score, which measures 12 aspects of physical functionality. It consists of 12 questions, each targeting a specific aspect of functional capability, with individual scores ranging from 4 (normal function) to 0 (complete loss of function). Thus, the aggregate score can span from a maximum of 48, indicating full functionality, to a minimum of 0, reflecting total loss of physical function. For a repeatedly measured continuous variable such as ALSFRS-R score, Mixed Models Repeated Measures (MMRM) analyses are typically employed.
A significant advantage of a platform trial is the comparison of patients receiving each active treatment to placebo patients shared from other treatment arms (regimens). The sharing of placebo patients (which we will call controls) from other treatment arms (regimens) allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, when employing models like MMRM to analyze repeatedly measured continuous variables such as the ALSFRS-R score, additional considerations emerge beyond the typical variability in estimated efficacy due to patient-to-patient differences within each treatment group. Concerns arise specifically regarding the influence of shared control samples. An overachieving shared control group could inadvertently suggest efficacy across all compounds in a platform trial, presenting challenges for sponsors. Conversely, an underperforming shared control group might lower the bar for all compounds to pass regulatory scrutiny too easily. This article addresses the necessity of planning for a sufficient total sample size to address these concerns while accommodating various randomization ratios that satisfy all stakeholders: patients, regulators, and sponsors.
2 False Approval Rate (FAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the purpose of Type I error rate control is to control the rate of approving a treatment (e.g., drug) when it is not efficacious, which we refer to as False Approval Rate (FAR) control. To balance patient benefits and risks, the US FDA requires pharmaceutical companies to provide evidence of a new drugs’ safety and efficacy before it is approved for public use, for which the FDA permits a maximum of five incorrect approvals for every 100 ineffective compounds tested, effectively a 5-per-100 rate. To achieve this, the FDA mandates controlling familywise error rate at $5\%$ within each study. In this paper, we refer to each clinical trail as a “study” and we will show that across studies, FAR is controlled, regardless of whether the studies are independent or correlated in families of platform trials.
2.1 Two layers of families and error rates
Within study
Within each study, there may be multiple doses and/or endpoints. The within-study family of inferential statements is the collection of (prespecified) statements to be made about them (e.g. confidence intervals for low dose efficacy in the primary endpoint and high dose efficacy in the secondary endpoint).
Across studies
Across all studies, the outcome of each study is whether the study compound is approved or not. The across-study family of inferential statements is the collection of statements to be made about whether the test compound of each study in the family has efficacy or not.
In general, when there are decisions to be made both within and across studies, multiple layers of families may exist. For instance, Ding et al., (2018) has a lower layer family comprising three confidence intervals for each single nucleotide polymorphism (SNP), assessing its efficacy across three possible genotypes. At a higher layer, there are confidence statements across all the SNPs in a genome-wide association study (GWAS), evaluating whether each SNP is linked to the phenotype, thus demonstrating the efficacy of a new treatment in comparison to a control. As another example, Han et al., (2022) has a lower layer family of three confidence intervals for each biomarker value, assessing the efficacy of a new treatment versus a control within three distinct groups: the marker-positive $g^{+}$ subgroup, marker-negative $g^{-}$ subgroup, the overall population $\{g^{+},g^{-}\}$. At a higher layer, there are confidence statements across uncountable many biomarker values.
In the Healey ALS Platform Trial, within each regimen, there exists a family of efficacy assessment on primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial form a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
Principle 2.1 (Principle of layering of error rate controls).
When there are multiple layers of families of statistical inferences, error rate control can be applied in layers, to control the incorrect decision rate at each layer as well as for the overall decision-making process.
2.2 Tukey’s three Multiple Comparison error rates
It can perhaps be said that Tukey (1953), which has been reprinted as Tukey (1994), started Multiple Comparisons as a discipline in Statistics. The form of statistical inference that John W. Tukey preferred was confidence interval. He defined the following error rates:
error rate per comparison
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of comparisons}}$
error rate per family
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of families}}$
error rate familywise
$=\frac{\mbox{number of erroneous families}}{\mbox{number of families}}$
In these definitions, each “statement” is a confidence interval, and an erroneous statement is a confidence interval that does not cover the true parameter value. An erroneous family is a family with at least one confidence interval not covering its true parameter value.
Note the now common phrasing familywise error rate transposes the error rate and familywise in Tukey’s original terminology. We believe the intent of the per in error rate per family was to have an indication of “how many?” after that. An error rate of 1 per a family of 20, or equivalently an error rate of 5 per a family of 100 (5-per-100 in short), both are $\mathbf{5\%}$ Tukey’s ordering of words is evocative because saying
  • error rate per a family of $k$ studies indicates control of errors is across studies
  • error rate familywise indicates control of errors is within each study
Therefore, instead of the transposed familywise error rate (abbreviated as FWER), we use Tukey’s original ordering error rate familywise (abbreviating it as ERFw). Also contrary to popular perception, it is not true that the only alternative to ERFw is false discovery rate (FDR). While FDR control may be appropriate for medical image reconstruction, it is well known that FDR control is inappropriate for controlling errors within a clinical trial because it can be manipulated by sponsors (see Section 6 of Finner and Roter (2001)). Which multiple comparison error rate to control depends on the specific application as we will discuss below.
2.3 Error Rate Familywise (ERFw) control Guards regulatory False Approval Rate (FAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to an incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to an incorrect claim of additional efficacy on the drug label. Since any incorrect statement can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise (ERFw) within each study since the early 2000s.
Different treatment arms within a platform trial may share a common control arm and therefore are not independent. Across the infinitely many studies in independent trials and in platform trials, the following theorem allows us to show if error rate familywise (ERFw) of every study within each platform trial is set at the 5% level, the aforementioned dependence does not affect regulatory false approval rate being no more than 5-per-100 (or 5%), so long as dependence is short-term (within each platform trial only, while studies in different platform trials remain independent).
Theorem 2.1 (Law of Large Numbers for random variables with short term dependence).
Let $X_{l},X_{2},\ldots,$ have means $\mu_{1},\mu_{2},\ldots$ variances $\sigma^{2}_{1},\sigma^{2}_{2},\ldots$ and covariance $Cov\{X_{i},X_{j}\}$ satisfying
$Cov\{X_{i},X_{j}\}\leq\rho_{j-i}\sigma_{i}\sigma_{j}~{}(i\leq j)$ (2.1)
where $0\leq\rho_{m}\leq 1$ for all $m=0,1,\ldots$. If the series $\sum_{1}^{\infty}\rho_{m}$ and $\sum_{1}^{\infty}\sigma^{2}_{i}(\log i)^{2}/i^{2}$ are both convergent, then as $n\rightarrow\infty$
${n}\sum_{i=1}^{n}X_{i}-\frac{1}{n}\sum_{i=1}^{n}\mu_{i}\rightarrow 0~{}{w.p.1}.$1 . (2.2)
Proof.
See Theorem E on page 27 of Serfling, (1980). ∎
Even though individual studies and those in platform trials are conducted at overlapping times, we can think of them as occurring in a (countably) infinite sequence of clinical trials, indexed as Study $\ldots,-1,0,1,\ldots$. Consider a finite number of $k$ studies, some of which are in platform trials while others are not. Within these $k$ studies, some compounds have efficacy while others do not. Studies with efficacious compounds that fail, i.e. failing to reject the null hypothesis of no-efficacy, do not lead to incorrect regulatory approval. We thus isolate those studies with compounds that lack efficacy (and therefore should not gain regulatory approval) as studies with index in $N$, and let $n$ denote its cardinality (the number of studies with index in $N$).
Let $\ldots,Y_{-1},Y_{0},Y_{1},\ldots$ be the indicator functions of past, present, and future studies that lead to efficacy claims from rejection of the null hypothesis that the compound studied lacks efficacy. Among studies with index in $N$ that should not lead to regulatory approvals, we relabel their $Y_{i}$ as $X_{i}$, i.e., $X_{i}$ is the indicator of a Type I error being committed as a result of Study $i,~{}i\in N,$ a study with compound that lacks efficacy,
$X_{i}=\left\{\begin{array}[]{ll}1&\mbox{if a Type I error is made as a result of Study}~{}i,i\in N,\\ 0&\mbox{otherwise}.\end{array}\right.$ (2.3)
Then $E(X_{i})=\mu_{i}$ for all $i\in N$. For $X_{i},X_{j}$ within the same platform trial, the correlation $\rho_{j-i}$ in (2.1) is between zero and one. But with short term dependence, for $X_{i},X_{j}$ in different platform trials, $\rho_{j-i}=0$. For example, if the practical maximum number of study arms within a single Platform trial is 100, then the $\rho_{j-i}$ in (2.1) would satisfy
$\rho_{j-i}\left\{\begin{array}[]{ll}\leq 1&\mbox{if }|j-i|\leq 100\\ =0&\mbox{if }|j-i|>100.\end{array}\right.$
Let $m$ = $|j-i|$, then $\rho_{m}=0,~{}m=101,102,\ldots$ for platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series.
Let $V$ denote the number of studies incurring Type I errors in $N$, i.e.,
$V=\sum_{i\in N}X_{i}.$
Assuming $P(X_{i}=1)=5\%$ for every $i\in N$, the expectation of $V$ can be expressed as the sum of the probabilities for each $X_{i}$, leading to:
$E[V]$ $=$ $\sum_{i\in N}P(X_{i}=1)=5\%\times n.$ (2.4)
This additive multiplicity adjustment (2.4) for controlling ERpF is an exact equality regardless of the dependency structure among the $X_{i}$ variables, whether they are independent, positively dependent, or negatively dependent. It should in no way be confused with the Bonferroni inequality multiplicity adjustment for the purpose of controlling ERFw.
Moreover, Theorem 2.1 assures that as the number of studies ($n$) grows infinitely large, the proportion of studies incurring Type I errors ($V/n$), i.e., ERpF, will converge to the probability of a Type I error in any single study (i.e., ERFw)—$5\%$ in this case. In other words, by the additive multiplicity adjustment (on line 360 of Food and Drug Administration, (2023)), controlling the Type I error rate familywise (ERFw) of each study exactly at 5% leads to controlling the error rate per family (ERpF) at exactly 5% for a family of $n$ studies with in-efficacious compounds (which we phrase as 5-per-100). More generally, regardless of whether studies in the families are independent or dependent, having an error rate per family no more than $5\%\times n$ for a family of $n$ studies means to expect no more than one incorrect approval per a family of 20 studies, five incorrect approvals per a family of 100 studies, etc. Such precision in error control exemplifies the regulatory commitment to balancing risk and benefit for patients, challenging the perception suggested by Bai et al (2020) Bai et al., (2020).
3 Stability of FAR and Conditional Assessment
While reassuring that a 5% ERFw within each study results in an ERpF of 5% across all the studies, there is a concern regarding the stability of ERpF across platform trials. Lines 365 and 366 in Section F, “Multiplicity”, of chapter III , “Considerations On Design and Analysis,” in the Food and Drug Administration (2023) guidelines, specifically state that
the probability distribution for the number of type I errors should be considered both in evaluating a proposed design and analysis plan and in evaluating the persuasiveness of results.
This section provides a clear demonstration of how to meet this requirement.
Note that it is not whether the unconditional or the conditional error rate should be controlled, as not one error rate control is more meaningful than the other. Both should be assessed because, from elementary statistics, only if one knows the marginal (unconditional) as well as the conditional can one sees the whole picture, i.e. the joint distribution. Thus, think of the platform-specific conditional error rate calculation we describe in this section as within the controlled unconditional error rate over the infinitely many studies over time, conditional on the specific number of studies for a specific platform trial (5 say), with a specific value for the sample mean of the shared control (one SE below its true value say).
3.1 Conditional Stability of FAR is desirable
An error rate represented as an expectation reflects the average value in the probability distribution of errors made. Although an unconditional error rate might be controlled at $5\%$, actual error rates can vary significantly among different families of studies, depending on the characteristics of studies within each family. For example, some families may experience error rates ranging broadly from $1\%$ to $9\%$, while others might have more stable rates, fluctuating only between $4\%$ and $6\%$. Thus, it’s important to consider not only the unconditional (expected) error rate but also the variability of the actual error rates, conditioned on the specific characteristics of the studies within each family.
Let $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy. For each $i\in N$, let $Z_{i}$ be the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)=\alpha$. Define $V^{*}$ as the sum of these indicators:
$V^{*}=\sum_{i\in N}Z_{i}.$
The variance of $V^{*}$, given by
$Var(V^{*})=\sum_{i}Var(Z_{i})+\sum_{i\neq j}Cov(Z_{i},Z_{j})$
where $Var(Z_{i})=\alpha(1-\alpha)$, can be calculated once the covariances $Cov(Z_{i},Z_{j})$ for $i\neq j$ are known. This section will discuss two specific concerns on variability of $V^{*}$ associated with sharing a control in a platform trial.
3.1.1 Effect of correlation induced by sharing control on variability of $V^{*}$
In platform trials, there are several methods to compute the covariances of non-coverage events across different treatment arms. Specifically, let $T^{\star}_{i}$ represent the pivoted statistic used to compare the new compound $i$, $i\in N$, against a shared control in a platform trial, such that
$T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$ (3.1)
has a null distribution free of the efficacy parameter, where $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are estimated mean of the new compound $i$ and the shared control, respectively. For proof of concept, consider the case where each pair $(T^{\star}_{i},T^{\star}_{j})$, with $i\neq j$, follows a bivariate standard Normal distribution. This scenario is particularly applicable in therapeutic areas like Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM), where the primary efficacy endpoints11footnotetext: which excludes the score test version of the logRank test are well-defined and standardized. The covariance between non-coverage indicators $Z_{i}$ and $Z_{j}$ is given by:
$Cov(Z_{i},Z_{j})=P\{T^{\star}_{i}>z_{\alpha}\cap T^{\star}_{j}>z_{\alpha}\}-\alpha^{2},$ (3.2)
which depends on $\rho_{ij}=Corr(T^{\star}_{i},T^{\star}_{j})=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$.
Each subset $N$ has its distinct $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio of sample sizes between the treatment arms with index in $N$ and the shared control arm, but also on other covariates like baseline measurement of each patient. To assess the stability of error rate control for compounds apparently with no efficacy, we consider the observed treatment outcomes as well as the covariates as known. Consequently, we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ among the compounds in $N$ that show no efficacy. Such computations are feasible within model-based analytical frameworks.
Table 1 illustrates that the standard deviation of $V^{\star}$, denoted as $StdDev(V^{\star})$, increases with higher correlation coefficients $\rho_{ij}$. If the treatment arms were independent, $StdDev(V^{\star})$ exceeds the mean $E(V^{\star})$. Moreover, correlations among the test statistics, caused by sharing a control, further increase the variability of $V^{\star}$ (the number of incorrect approvals). For example, a common correlation of 0.5 among the test statistics raises $StdDev(V^{\star})$ from 0.69 to 1.16. This inherent variability of the actual error rate in a specific family of clinical trials is a significant issue that, we believe, warrants greater attention in practice, irrespective of whether the error rate is measured as ERpF or ERFw.
$k$ $E(V^{\star}$ ) $StdDev(V^{\star})$
Unconditional Independent studies Hypothetical Platform $k:1$ randomization
$n^{\mathrm{true}}$ ERpF $\rho_{ij}=0$ $\rho_{ij}=0.3$ $\rho_{ij}=0.5$
5 0.25 0.49 0.57 0.66
10 0.5 0.69 0.94 1.16
20 1 0.97 1.65 2.15
40 2 1.38 3.02 4.12
Table 1: For a family of studies, if $StdDev(V^{\star})>E(V^{\star})$, that is, standard deviation of the number of errors is larger than the mean of the number of errors, then one might consider stability of the ERpF control to be of concern. In a Platform trial, pairs of $T^{\star}_{i}$ and $T^{\star}_{j}$ are correlated due to sharing of Control patients. Assuming $\rho_{ij}$ is common for all pairs, variability of $V^{\star}$ increases with increasing $\rho_{ij}$. For illustration, consider a Platform trial with $k$ new compounds using a $k:1$ compound-to-placebo randomization (which makes the sample size for comparing each new compound with the shared control $1:1$). With ERpF set at $5\%\times n^{\mathrm{true}}$ where $n^{\mathrm{true}}$ is the number of in-efficacious compounds, we see independent studies reach stability by family size 20, but that is not the case for studies in platform trials. Stability of ERpF control is thus more of a concern in platform trials.
To facilitate the assessment of the impact of correlations, analogous to how internet commerce tracks cookies, we propose maintaining an informational “cookie” for each treatment or regimen in a platform trial.
Control Sharing Matrix
Each treatment or regimen should have a “cookie” that records the exact time interval of its enrollment period. By identifying the overlap between the enrollment periods of different regimens, it becomes possible to compute the pairwise covariance and correlation for comparisons between the active treatment and the control.
$n$ A B C D
A 122 122 164 103
B 122 120 122 103
C 164 122 120 103
D 103 103 103 75
Table 2: Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial
A B C D
A 1 0.498 0.425 0.478
B 0.498 1 0.496 0.476
C 0.425 0.496 1 0.476
D 0.478 0.476 0.476 1
Table 3: Correlations matrix between $\hat{\mu}_{i}-\hat{\mu}_{SC_{ij}}$ and $\hat{\mu}_{j}-\hat{\mu}_{SC_{ij}}$
for the HEALEY ALS Platform Trial
A B C D
A 0.0475 0.0093 0.0083 0.0087
B 0.0093 0.0475 0.0092 0.0096
C 0.0083 0.0092 0.0475 0.0086
D 0.0087 0.0096 0.0086 0.0475
Table 4: Matrix of variances of $Z_{i}$ (diagonal) and covariances between $Z_{i}$ and $Z_{j}$ (off-diagonal) with ERpF set at $5\%\times 4=0.2$ for Regimens ABCD of the ALS Platform trial.
Table 2 presents the Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial, inferred from the results available on clinicaltrials.gov. For each regimen, the diagonal elements represent the number of patients randomized to the active treatment. The off-diagonal elements indicate the number of patients shared in the control group, to which each active treatment is compared, for each pair of regimens. We can then derive Table 3, which presents the correlation matrix (Details of the derivation are available in the supplementary materials). Note that as mentioned before, the pairwise correlation in Table 3 can also be expressed as $\hat{\rho}_{ij}=\lambda_{i}\lambda_{j}$ for all $i<j\in N$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$. These $\lambda$ values can be readily calculated using existing software, making this notation preferable for its ease of extension to more complex structures. For the HEALEY ALS Platform Trial, the approximated $\lambda=$ (0.674, 0.722, 0.671, 0.690). Based on equation (3.2), we can obtain variance-covariance matrix for $Z$ (see Table 4). Note that the elements sum to 0.2974 for $Var(V^{\star})$, resulting in $StdDev(V^{\star})$ = 0.5454, which exceeds the ERpF of 0.2.
3.1.2 Effect of sample mean of shared control on variability of $V^{*}$
Assuming, without loss of generality, that $\mu_{i}=0$ for $i\in N$. According to equation (3.1), a positive (or negative) estimated mean $\hat{\mu}_{SC}$ of the shared control can make it harder (or easier) for new compounds to obtain approval. In this paper, a positive $\hat{\mu}_{SC}$ is referred to as an under-performing estimated mean of the shared control, while a negative $\hat{\mu}_{SC}$ is termed an over-achieving estimated mean of the shared control.
An under-performing $\hat{\mu}_{SC}$ presents significant concerns for sponsors seeking treatment approvals. The higher $\hat{\mu}_{SC}$ becomes, the more challenging it is for treatment arms to receive approval. This scenario leads to high probability of $V^{\star}$ being towards lower numbers of erroneous approvals, resulting in a lower variability of $V^{\star}$. However, for regulators, especially when the treatment arms are non-efficacious, this situation is less problematic as it reduces the likelihood of erroneously approving a treatment.
With over-achieving $\hat{\mu}_{SC}$, the situation becomes a concern for regulators, as it is much easier to approve treatment arms due to the lower $\hat{\mu}_{SC}$. This increased ease of approval heightens the risk of erroneous approvals. In extreme cases, if $\hat{\mu}_{SC}$ is sufficiently low, almost all treatment arms may be erroneously approved. While sponsors might favor such an outcome, it is undesirable from a consumer and regulatory perspective because it leads to the approval of non-efficacious drugs, thereby increasing the probability of higher $V^{\star}$ and therefore enhancing variability of $V^{\star}$.
These two cases are demonstrated in Figure (1). Assume $\mu_{i}=0$ for $i=1,2,3,4\in N$, and $\alpha=0.05$. Then, Figure (1) compares the true probability distribution of $V^{*}$ when $\hat{\mu}_{SC}={\mu}_{SC}=0$ (the dash line) with those where $\hat{\mu}_{SC}$ is either greater than or less than ${\mu}_{SC}=0$ (the solid line). Figure 1a indicates an essentially negligible change in the probability of obtaining $V^{*}=1$, and the probabilities of obtaining $V^{*}=2,3,4$ are nearly zero, suggesting minimal impact on the variability of $V^{*}$. However, as shown in Figures 1b and 1c, the likelihood of obtaining $V^{*}=1$ and $V^{*}=2$ increases as $\hat{\mu}_{SC}$ decreases, thereby increasing the variability of $V^{*}$.
(a) $\hat{\mu}_{SC}=0.46$
1 SE above $\mu_{SC}$
(b) $\hat{\mu}_{SC}=-0.23$
0.5 SE below $\mu_{SC}$
(c) $\hat{\mu}_{SC}=-0.46$
1 SE below $\mu_{SC}$
Figure 1: Probability distribution of $V^{\star}$. In this scenario, we have 5 treatment arms, each with an $\alpha$ level of 0.05, resulting in an ERpF of $0.05\times 5=0.25$. The total sample size is 1200, with a 5:1 ratio of treatment arm to control sample size allocation, leading to an unconditional correlation of 0.5 between estimates in differences between each new treatment arm and the control. For illustration purpose, we assume that the true shared control mean $\mu_{SC}$ and treatment arm means $\mu_{i}$ both equal to 0 (although these values can be adjusted). In Figure 1(a), the distribution of $V^{\star}$ closely mirrors the underlying true distribution. With a higher $\hat{\mu}_{SC}$, it becomes more difficult for treatment arms to gain approval, leading to a negligible probability of $V^{\star}>0$. Conversely, Figure 1(b) and 1(c) show scenarios where the $\hat{\mu}_{SC}$ is much lower than $\mu_{i}$, facilitating easier erroneous approvals. In particular, a substantial increase for $V^{\star}=1$ and $V^{\star}=2$ would not cause only an increase in $V^{\star}$, but also its variability.
3.1.3 Effect of mean and correlation induced by sharing control on variability of $V^{*}$
It is to be noted that while both the mean and the correlation induced by sharing control individually affect the distribution of $V^{\star}$, it is important to consider their combined effect on the distribution of $V^{\star}$. Specifically, the variance of $V^{\star}$ is noteworthy. We standardized $\hat{\mu}_{SC}$ by $SE_{SC}$ which we denote by $\hat{\mu}_{SC}^{\star}$, i.e.
$\hat{\mu}_{SC}^{\star}=\frac{\hat{\mu}_{SC}}{SE_{SC}}.$
As seen in Figure 2, for cases when $\hat{\mu}_{SC}^{\star}$ is greater than the treatment arm mean $\hat{\mu}_{i}$, the performance of the standard deviation of $V$ is lower and much more stable, which translates to a more stable $V^{\star}$ value. On the other hand, when $\hat{\mu}_{SC}^{\star}$ is less than $\hat{\mu}_{i}$, the behavior of the standard deviation of $V$ is higher and much more variable, which translates to a more variable $V^{\star}$. Due to this behavior of $V^{\star}$, it is beneficial to look at the whole behavior of the probability distribution of $V$ rather than single instances of $V$.
Figure 2: The behavior of standard deviation for $V^{\star}$ is heavily affected by correlation and $\hat{\mu}_{SC}$. Assume $\mu_{SC}=0$, i.e. it is non efficacious. As we decrease the value of our $\hat{\mu}_{SC}$ and thus $\hat{\mu}_{SC}^{\star}$, the performance of standard deviation of $V^{\star}$ becomes less stable.
$\hat{\mu}_{SC}^{\star}$ = -3
$\hat{\mu}_{SC}^{\star}$ = -1.5
$\hat{\mu}_{SC}^{\star}$ = 0
$\hat{\mu}_{SC}^{\star}$ = 1.5
$\hat{\mu}_{SC}^{\star}$ = 3
Figure 3: Snapshot of different standardized $\hat{\mu}_{SC}$ values on correlation’s effect on standard deviation where $\mu_{i}=0$. As can be seen, when the $\hat{\mu}_{SC}^{\star}$ is less than the treatment mean, the behavior is less simple whereas for $\hat{\mu}_{SC}^{\star}$ is greater than the treatment mean, the standard deviation of V decreases as correlation increases.
3.2 Conditional assessment of FAR control
A foundational principle in (frequentist) Statistics is
Principle 3.1 (The Conditionality Principle).
If, among many experiments that can inform on an unknown parameter, one is randomly selected and performed, then the information available about the unknown parameter depends only on the experiment actually performed.
See J. O. Berger (1985) page 31 and Casella and R. L. Berger (1990) page 268 for details.
To motivate the application of this principle to conditional confidence, consider Fieller’s confidence interval for the ratio of the means of two Normal distributions (Fieller 1954) which unconditionally (over infinitely many replications of a study) has confidence level exactly equal to $(1-\alpha)$. Interestingly, depending on the data observed in each particular study, Fieller’s confidence interval can be an interval, or the entire real line, the complement of an interval, or the empty set (see Casella and Berger 1990, page 459). To us, conditional on Fieller’s interval being the entire real line for a particular study, a conditional confidence of 100% should be stated for that study. Likewise, conditional on Fieller’s interval being the empty set for a particular study, a conditional confidence of 0% should be stated for that study, not averaging the coverage probability over studies that have not been performed that would result in Fieller’s confidence interval being an interval, the complement of one, or the entire real line. In this spirit, while taking comfort that a 5% ERFw within each study results in an unconditional ERpF of 5% across all the studies, we recommend assessing the conditional ERpF within each Platform trial conditional on the observed shared Control outcome, to be sure they do not deviate too much from the unconditional ERpF.
4 A Confident Platform Random Sizer App
An app [^2] Can be found at https://eouya001.shinyapps.io/ConfidentPlatformRandomSizer/ has been developed to not only illustrate the impact of conditioning on $\hat{\mu}_{SC}$ for the distribution of $V^{*}$, but also to provide insights into the calculated standard errors and unconditional correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. We simulate a case study where we test the efficacy of five drugs. The total resources available allow for 600 people. With this setup, we can adjust the input values using the app’s sliders. Notably, ALS has a population standard deviation of 6.5 ($\sigma$=6.5), which the app uses internally to calculate standard errors of $\hat{\mu}_{SC},\hat{\mu}_{i}$ and $\hat{\mu}_{i}-\hat{\mu}_{SC}$, $i=1,\ldots,n\in N$, where $n$ is the maximum number of treatment arms erroneously approved. An example case in which sponsors should be concerned is presented in App: User Slider Controls.
App: User Slider Controls $\alpha$ for Each Rx Arm: ERFw $\alpha$ for each treatment arm. Note that ERpF is $n\alpha$
$\alpha=0.05$ for each treatment arm.
Standardized $\hat{\mu}_{SC}$: The shared control sample mean $\hat{\mu}_{SC}$, standardized by the standard error of the shared control sample mean
An example of an overachieving $\hat{\mu}_{SC}$, i.e. $\hat{\mu}_{SC}>0$. Different shared control values would affect the number of erroneous approvals
Maximum Number of Rx Arms Erroneously Approved: $n$, the number of treatment arms, when they are non-efficacious.
n=5
Total Sample Size: Sum of sample sizes from all treatment arms and the shared control
Total sample size = 600
Unconditional Correlation Between New Treatment and Control Estimates: Correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
$\rho_{ij}=0.5$
[Uncaptioned image] Sample Size Ratio: Treatment to control sample size ratio for sample allocation. Note that while the app allows for one to go higher, it is generally not advised to go higher than $k:1$.
5:1 treatment-to-control sample size ratio
By selecting the sample size ratio, the app automatically calculates the unconditional correlation between the new treatment arm and control estimates. This feature is particularly beneficial to regulators, who have a vested interest in managing this aspect to stabilize the distribution of $V^{*}$. A lower correlation achieved by reducing the sample size ratio also decreases the variation of $V^{*}$.
App: Output Distribution of $V^{*}$: Gives the expected value, standard deviation, and the probability distribution of $V^{*}$, the number of treatment arms erroneously approved.
With an overachieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
Sample Allocation: Gives the recommended sample sizes for each treatment arm and the shared control (rounded to nearest whole number)
The study utilizes as 500:100 ratio, so a ratio of 5.
[Uncaptioned image] Standard Errors: Gives standard errors for estimated control mean, treatment arm means, and the mean difference between control and treatment arm
Standard errors were calculated based on the standard deviation $\sigma$ = 6.5
The app can be instrumental in designing new studies. It facilitates considerations regarding the stability of the estimates from shared controls and the (unconditional) correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. Specifically, if regulators set a required correlation threshold, sponsors can use the app to determine the maximum sample ratio randomization that meets this criterion. The app thus serves as a valuable tool for identifying the optimal balance that satisfies all stakeholders involved.
Within a platform trial, the randomization ratio $n_{Rx}:n_{C}$ of the new compound to the shared control can vary. Ratios range from a low of $1:1$ resulting in a sample size ratio of the shared control to each new compound of $k:1$—to a high of $k:1$, where the sample size ratio reverses to $1:1$. Given a fixed total sample size for the platform trial, selecting the appropriate randomization ratio involves considering multiple factors.
Patients generally prefer a higher $n_{Rx}:n_{C}$ ratio, as it increases their chances of receiving a new treatment compound. For individual drug sponsors, the advantage of participating in a platform trial includes being able to pool control patients, which allows for a higher randomization ratio than the standard $1:1$ typically seen in standalone studies. This higher ratio can facilitate more accurate effect estimations for new compounds.
However, the perspective of the master protocol sponsor tends to favor a lower ratio, which increases the sample size of the shared control, reducing the likelihood of an overachieving control sample that complicates the demonstration of efficacy across new compounds.
From a regulatory standpoint, the stability of the ERpF control is influenced by both the estimates of the shared control’s sample means and the correlation among the pivotal statistics $T^{*}$. A lower randomization ratio enhances the sample size of the shared control, mitigating the risk of an under-performing control sample that could erroneously ease the approval of ineffective compounds. Additionally, a lower ratio reduces the correlation among pivotal statistics, thereby enhancing the stability of ERpF control.
For example, in the HEALEY ALS Platform Trial, none of the four drugs tested in Regimens ABCD demonstrated efficacy in the primary endpoint, ALSFRS-R. [^3] Active treatments for Regimen A were Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine. The lack of efficacy could be attributed to either the inherent ineffectiveness of the compounds, an overachieving shared control, or a combination of both. Conversely, as illustrated in App: Regulator Concerns, with an under-performed shared control sample mean had (i.e., $\hat{\mu}_{SC}<0$), there might have been a higher chance of the drugs being approved too readily, as exemplified by a $\hat{\mu}_{SC}^{\star}=-1.5$. Such a scenario would raise significant concerns for regulators.
In sum, choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these considerations into account. Our app lets one find the sweet spot.
App: Regulators Concerns We now present a scenario using our app where sponsors, such as MASS General, need not be concerned, but which should capture the attention of regulators. The following inputs and outputs illustrate a hypothetical situation: $\alpha$ for Each Rx Arm: $0.05$ Standardized $\hat{\mu}_{SC}$: $-1.5$ Maximum Number of Rx Arms Erroneously Approved: $n=5$ Total Sample Size: 600 Unconditional Correlation Between New Treatments and Control estimates: 0.5 Sample Size Ratio: 5:1
[Uncaptioned image] The probability distribution of $V^{*}$ indicates that an under-performing standardized $\hat{\mu}_{SC}$ significantly increases the likelihood of erroneously approving at least one treatment, compared to its true efficacy. Regulators should be particularly concerned about this, as it suggests a non-negligible probability that ineffective drugs could mistakenly reach the market.
5 False Discovery Rate (FDR) in a regulatory setting
5.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed as a method for controlling the Type I error rate across studies in Platform trials (Robertson et al., (2023)). FDR is defined as the expected proportion of incorrect rejections (rejecting a true null hypothesis) among all rejections, both correct and incorrect  [^4] Rejecting a false null hypothesis is considered “correct”, while rejecting a true null hypothesis is “incorrect”..
Let $H_{0i},i=1,\ldots,k$ represent the null hypotheses that each compound in a Platform trial with $k$ studies lacks efficacy, then
$\mathrm{FDR}$ $=$ $\mathrm{E}\left[\frac{\mbox{number of rejected }H_{0i}\mbox{ thatare true}}{\mbox{total number of rejected }H_{0i}}\right]$
$=$ $\mathrm{E}\left[\frac{\mbox{number of in-efficacious compounds incorrectly approved }}{\mbox{total number of compounds approved, both efficacious and in-efficacious}}\right]$
While current regulatory practice of controlling FAR (which is Tukey’s per family error rate across studies) limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100 as we will show below.
Recall $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy and $|N|=n$. For each $i\in N$, $Z_{i}$ is the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)$, and
$V^{*}=\sum_{i\in N}Z_{i}.$
Then
$\dfrac{\mathrm{EFpF}}{n}$ $=$ $\dfrac{E(V^{*})}{n}$ (5.2)
$=$ $\dfrac{\sum_{i=1}^{n}P\left\{Z_{i}=1\right\}}{n}$ (5.3)
$=$ $\sum_{v=1}^{n}\dfrac{v}{n}P\{V^{*}=v\}$ (5.4)
Now suppose a Platform trial has $n^{\mathrm{false}}$ extremely false no-efficacy nulls (that will almost surely get rejected), then
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{n}\dfrac{v}{v+n^{\mathrm{false}}}P\{V^{*}=v\}$ (5.5)
$\leq$ $\sum_{v=1}^{n}\dfrac{v}{0+n^{\mathrm{false}}}P\{V^{*}=v\}.$ (5.6)
How ERpF and FDR count the number of incorrect approvals $v$ differently is evident by comparing (5.5) with (5.4). FDR attenuates the number of incorrect rejection $v$ by the number of correct rejection of false nulls which we denote by $R$.
From (5.6), we see $n^{\mathrm{false}}$ being not small (relative to $v$) allows inflation of $P\left\{Z_{i}=1\right\}$ while keeping FDR small. For simplicity, suppose each Platform trial has $k=2$ studies only, and suppose the compound in Study $w$ is highly efficacious but the compound in Study $v$ is in-efficacious:
  • $H_{0w}^{{\scriptsize\mbox{false}}}$ is so false that $P\left\{\mbox{Reject }H_{0w}^{{\scriptsize\mbox{false}}}\right\}=1$;
  • $H_{0v}^{{\scriptsize\mbox{true}}}$ is true so rejecting it constitutes the only incorrect rejection.
In this scenario, $H_{0v}^{{\scriptsize\mbox{true}}}$ can be tested at a Type I error rate (ERFw) of $10\%$ (within Study $v$) while keeping FDR in this Platform trial of a family of 2 studies at $5\%$, because
$FDR$ $=$ $E\left[\frac{V}{R}\right]$
$=$ $\left(\dfrac{1}{1+1}\right)\times P\left\{\mbox{Reject }H_{0v}^{{\scriptsize\mbox{true}}}\right\}$
$=$ $\frac{1}{2}\times 10\%=5\%$
So, if there is a sequence of many Platform trials with $k=2$ studies following this scenario, then by the LLN (2.1), the False Approval Rate (FAR) is 10-per-100 (not 5-per-100).
Switching to FDR control would thus be a policy change: For every efficacious compound approved, regulators reward the industry with an additional fungible False Approval Rate (FAR) token [^5] We term the tokens fungible because any company can cash them in for any compound. toward approval of in-efficacious compounds.
Adopting FDR control would represent a significant policy shift. Under this system, for every efficacious compound approved, regulators would essentially reward the industry with an additional fungible False Approval Rate (FAR) token [^6] We describe the tokens as fungible because any company can use them to seek approval for any compound. that could be used toward the approval of inefficacious compounds. It is important to note that we find no explicit language in the Food, Drug, &\&& Cosmetic Act or the Kefauver-Harris Amendment suggesting that the number of efficacious compounds correctly approved affects the guarding of the number of in-efficacious compounds incorrectly approved.
5.2 Integrity of error rates against manipulation is desirable
Whereas the False Approval Rate (FAR) is only concerned with the rate at which in-efficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter (2001), we wonder whether that might subject FDR to manipulation as follows.
If FDR control were implemented across platform trials involving multiple drug companies. It seems plausible that a strategic company might hesitate to submit their best compound, preferring instead to let other companies take that risk. Similarly, in platform trials with a single sponsor, there could be a temptation to include highly efficacious compounds that have low commercial value. This strategy could potentially be used to facilitate the approval of less efficacious but more profitable drugs.
6 Conclusion
In this paper, we examined the error rates within clinical studies, particularly focusing on platform trials. Our findings offer significant insights into the control of error rates and their implications across various study designs, especially in the context of regulatory and clinical settings.
We demonstrated that controlling the ERFw within a single study at a threshold of $5\%$ effectively controls the ERpF across multiple studies at a rate of 5-per-100. This finding holds irrespective of the correlation between studies, including those grouped within platform trials. Furthermore, our analysis supports that current regulatory practices, which involve controlling the ERFw within a study, limit the False Approval Rate (FAR) to 5-per-100. This approach ensures that no more than five ineffective drug compounds receive approval per hundred evaluated, thus safeguarding public health while fostering innovation.
A significant advantage of a Platform trial is that multiple treatments are evaluated simultaneously against a shared control, which allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, the stability of regulatory FAR across Platform trials is of serious concern since shared controls introduce dependencies among treatment evaluations and therefore increases the variability of ERpf. As we demonstrated in our analysis, dependencies among treatment evaluations increases with the treatment/control randomization ratio. From patients perspective, a higher treatment/control randomization ratio will allow higher chance of receiving a new treatment compound. From an individual drug sponsor’s point of view, a higher treatment/control randomization ratio through participating platform trial allows more accurate estimate on the effects of new compounds. However, from the master protocol sponsor’s point of view, lower the the treatment/control randomization ratio allows increasing the sample size of the shared control, and therefore lessens the chance that an overachieving shared control sample causes all the new compounds have difficulty to demonstrate efficacy. Additionally, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable. To balance the different perspective from different stake holder, our Confident RandomSizer app offers a tool to identify optimal treatment/control randomization ratio given a fixed sample size.
In conclusion, maintaining stringent control of error rates within clinical trials is crucial for ensuring the reliability of study outcomes and the safety of approved treatments. As clinical trial designs become increasingly complex, particularly with the rise of platform trials, the statistical community and regulatory bodies must continuously adapt and refine their methodologies to address these challenges effectively.
7 Glossary
7.1 Terms
  • ALS: Amyotrophic lateral sclerosis (ALS)/Lou Gehrig’s disease; a fatal progressive neuromuscular disease
  • ERpF: Error Rate per Family; measures errors across studies
  • ERFw: Error Rate Family wise; error within each study
  • FAR: False Approval Rate; proportion of drug compounds that lack efficacy incorrectly approved
  • FDR: False Discovery Rate; expected proportion of rejections that are incorrect
  • FWER: Family Wise Error Rate; probability of type I error in a family
7.2 Notation
  • $\mu_{SC}$: Shared control sample mean
  • $\hat{\mu}_{SC}$: Shared control sample mean estimate
  • $\hat{\mu}_{SC}^{\star}$: Standardized shared control sample mean estimate
  • $X$: Indicator that type I error is committed as a result of $i$th study
  • $V$: The number of Type I errors committed in a family
  • $Z$: Indicator that $i$th confidence interval erroneously shows efficacy
  • $V^{\star}$: The number of treatments with their confidence intervals for efficacy not covering their true parameter values given $n$ confidence intervals
  • $SE$: Standard error
    • Subscripts denote standard error of respective term
  • $k$: Number of studies
  • $n$: Maximum number of treatment arms erroneously approved
  • $T^{\star}_{i}$: Statistic for comparing $k$ new compounds against shared control
    $T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$
  • $R$: the number of correct rejection of false nulls
  • $n_{SC}$: sample size for shared control
  • $n_{i}$: sample size for $i$th treatment
  • $\lambda_{i}$: Have that
    $\lambda_{i}=\left(1+\frac{n_{SC}}{n_{i}}\right)^{-\frac{1}{2}}$
  • $\rho_{ij}$: Correlation between sample statistics $T_{i}$ and $T_{j}$
    • Note: Can be written as a product of $\lambda_{i}$ and $\lambda_{j}$.